### v2

@@ A A D D B B C C A B C D A B C D A B C D
Table 4: Bases of the 600-cell. The generators $a$ through $e$, shown in bold font, give rise to the orbits of bases shown below them. The 75 bases in this table make up the basis table of the 600-cell. Note that the ray numbers keep increasing by one as one moves down a column, except when the end of a pentadecagon is reached, when one goes back to the ray at the beginning of that pentadecagon (”wraparound”). The bases associated with the generators $a$ and $b$ (which have the profiles AADD and BBCC, respectively) give rise to $30_{2}-15_{4}$ parity proofs involving thirty rays that each occur twice over fifteen bases.
Table 4 shows the orbits of fifteen bases arising from each of the generators $a$ through $e$ of Table 3. The 75 bases in Table 4 constitute the entire basis table of the 600-cell, which can be described by the word abcdeabcde. The bases associated with generators $a$ and $b$ each yield a $30_{2}-15_{4}$ parity proof, since they involve an odd number of bases (fifteen) with each ray occurring twice over them. It is interesting that the triacontagonal projection yields these proofs simply as a consequence of the way the bases are generated with it.
Any fifteen-fold symmetric parity proof of the 600-cell must be made up of a combination of the columns in Table 4. Accordingly, it can be written as a “word” made up of an odd number of distinct letters chosen from $a,b,c,d$ and $e$. An odd number of letters is needed to ensure that the total number of bases associated with the proof is odd. Additionally, the number of times each of the pentadecagons occurs over the letters of the word must be even to ensure that each ray occurs an even number of times over the bases. These two conditions restrict the allowed proofs to the words $a,b,acd,ace,ade,bcd,bce$ and bdebde. Let us term a parity proof “minimal” if it cannot be reduced to a shorter proof by leaving out some of the letters in it. Then the only minimal proofs in the preceding list are $a$ and $b$ because all the other proofs can be reduced to them by dropping two of their letters.
Although the triacontagonal projection yields just two fifteen-fold symmetric parity proofs of the 600-cell, it has billions of other proofs of the BKS theorem among its bases, as discussed in Waegell2011b and Lisonek2014b .
@@ 4 The 120-cell
The 120-cell is a four-dimensional regular polytope with 600 vertices lying on the surface of a 3-sphere. It has 300 rays that form 675 bases, with each ray occurring in nine bases, so its ray-basis symbol is $300_{9}-675_{4}$.
The triacontagonal projection arranges the rays at the vertices of twenty concentric regular pentadecagons, sixteen of which lie in pairs on eight different circles and four of which lie on circles of their own. Following ChiltonChilton , we label the pentadecagons by the letters A through L in order of decreasing circumradius, but add subscripts to eight of the letters to distinguish between pentadecagons lying on circles of the same radius; see Table 5 for the numbering of the rays and their coordinates in the plane of the projection.
The 675 bases of the 120-cell fall into forty five orbits of fifteen bases each. Table 6 lists one generator for each orbit, along with their basis profiles. Most profiles have only one orbit associated with them, but two of them have three each. Four of the generators, namely, those labeled $j,q,r^{\prime}$ and $s^{\prime}$, directly give $30_{2}-15_{4}$ parity proofs. This can be seen simply from their basis profiles, listed to their left, since each has the rays of two pentadecagons occurring twice over its fifteen bases. However there are many other parity proofs that can be constructed by combining the generators in Table 6, and we now discuss how that can be done.
Pentadecagon Circumradius Angle Rays
A 1.0000 6 1-15
B1 0.9515 1.76 16-30
B2 0.9515 10.24 31-45
C 0.9004 0 46-60
D1 0.8673 3.13 61-75
D2 0.8673 8.87 76-90
E1 0.8110 2.07 91-105
E2 0.8110 9.93 106-120
F1 0.7741 4.24 121-135
F2 0.7741 7.76 136-150
G1 0.6402 4.24 151-165
G2 0.6402 7.76 166-180
H1 0.5927 4.58 181-195
H2 0.5927 7.42 196-210
I1 0.5067 3.31 211-225
I2 0.5067 8.69 226-240
J 0.4452 0 241-255
K1 0.3219 1.76 256-270
K2 0.3219 10.24 271-285
L 0.0947 6 286-300
Table 5: Pentadecagons of the 120-cell, labeled A through L in order of decreasing circumradius, with pairs having the same circumradius being distinguished by a subscript. The rays are numbered in counter-clockwise order around the pentadecagons, beginning from the outermost pentadecagon and proceeding inwards. The ray with the lowest number on each pentadecagon is the one lying on or just above the positive x-axis. The starting angles of rays on all pentadecagons are indicated in the third column. The rays on any pentadecagon are spaced $24{{}^{\circ}}$ apart.
@@
Table 6: Generators of the 120-cell. The generators are shown in the second and fourth columns, along with their letter labels. The generators have the 41 basis profiles shown in the first and third columns, with the letters of any profile indicating the pentadecagons to which the corresponding rays of the generator(s) belong. Most of the profiles have just one generator associated with them, but two of them have three each. Each generator gives rise to an orbit of fifteen bases, and the union of all the orbits yields the 675 bases of the 120-cell.
We use a technique due to Lisonek, Raussendorf and Singhlisonek1 . Consider the matrix $M$ whose $i,j$-th element, $M_{i,j}$, is the number of times pentadecagon $i$ of Table 5 occurs in the basis profile of generator $j$ of Table 6 (we take $i$ and $j$ to be the order in which the pentadecagons or generators occur in Table 5 and 6). The matrix $M$ obtained in this way is a 20 x 45 matrix. A parity proof can be extracted from $M$ by picking an odd number of its columns in such a way that if they are stacked next to each other, the sum of the numbers in each of the rows is even. The word describing the resulting parity proof has for its letters the generators corresponding to the columns of $M$ that are picked. It remains only to explain how the right columns of $M$ are to be picked.
The problem is solved if we find all 45-dimensional column vectors $X$, with elements of 0 or 1, for which the equation $MX=0$ (mod 2) is satisfied, for then the non-vanishing elements of $X$ will indicate the columns of $M$ to be picked. The solutions to the foregoing equation are all the vectors in the nullspace of $M$ (by which we mean all vectors $X$ such that $MX=0$ (mod 2)). We used Maple to find that the nullspace has dimension 30 and obtained a set of linearly independent vectors in it. Taking all possible linear combinations of these vectors, with coefficients of 0 or 1, gives $2^{30}$ solutions to the equation $MX=0$ (mod 2). However, only half of these solutions (namely, those with an odd number of 1’s as elements) give parity proofs. Thus the number of parity proofs is $2^{29}$, and each of them can be constructed as the union of the bases associated with the generators picked out by the non-vanishing elements of $X$.
@@
Table 7: The 30 words representing a set of linearly independent vectors in the nullspace of the matrix $M$. The first word is of odd length, while all the others are of even length.
@@ 13 $bcdefghiknszq^{\prime}$ $225_{2}30_{4}15_{6}15_{8}$-$195_{4}$ (irreducible)
Table 8: Parity proofs of the 120-cell with all word lengths from 1 to 15 obtained by combining words in Table 7. Proofs with the same ray-basis symbol are listed together, with the symbol being shown in the last column. If the proof is reducible, the smaller proofs within it are indicated (see the text for an explanation of this point). The symbol $\oplus$ indicates that the smaller proofs have no bases in common, whereas + (which occurs only in the last entry) indicates that they do. See Table 9 for a blown-up version of the proof cdycdy. The dots … in the last column indicate that the ray-basis symbol to their left has to be added to itself the number of times shown in brackets.
@@ A C J L A D1 I1 L C D1 I1 J
Table 9: The three-letter proof cdycdy. The generators $c$,$d$ and $y$ are shown in bold font at the top and their orbits of fifteen bases below them. The entire set of bases in this table gives a $90_{2}-45_{4}$ parity proof, but the proof is reducible and breaks up into the three smaller $30_{2}-15_{4}$ proofs which are indicated in bold, italic and regular font. Each of the smaller proofs has a five-fold symmetry (i.e., adding three to the rays of any basis yields the next basis in it). The smaller proofs have no bases in common, so cdycdy is written as a direct sum ($\oplus)$ of them in Table 8.
The procedure just described can be translated into a simple algorithm for constructing the parity proofs. Let a “word” be a sequence of distinct letters representing the generators in Table 6. Then the 30 linearly independent vectors in the nullspace of $M$, which each pick out a set of generators, can be written in the form of the words shown in the Table 7. The words are listed in order of increasing length and the letters of any word are arranged in alphabetical order, with all primed letters coming after the unprimed ones (this convention eliminates the confusion caused by writing the same word in different forms).
The algorithm for obtaining parity proofs can now be stated: combine the words in Table 7 in all possible ways that lead to words of odd length. Each such word represents a parity proof.
To complete the description of the algorithm, we must explain how two or more words in Table 7 are to be combined. We give a procedure for combining two words which can be used repeatedly to combine any number of words from Table 7, even when one or both of the words do not belong to that table. Let $U$ and $V$ be the sets of letters corresponding to the two words that must be combined. Then the word obtained by combining them is represented by the set of letters $W$ that is the symmetric difference of the sets $U$ and $V$:
@@
Stated in words (pun not intended!), the word obtained by combining two words is the union of their sets of letters, but with all letters common to the two sets dropped. This prescription can be used to combine any number of words in Table 7. If the word that results is of odd length it represents a parity proof, otherwise it must be dropped. It is not hard to see that the words in the nullspace of $M$ form a group under the law of composition given in Eq.(1), with the null word being the identity element and every word being its own inverse.
To ease the task of getting directly at the odd length words, all but one of the words in Table 7 have been chosen to be of even length. This makes it possible to construct odd length words by combining any number of even length words to get an even length word and finally combining it with $j$ to get an odd length word.
An example may help make the construction clear. If one combines the 6-letter word abdekrabdekr with the 4-letter word $dgji^{\prime}$, one gets the 8-letter word $abegjkri^{\prime}$ and combining it with $j$ gives the 7-letter word $abegkri^{\prime}$, which represents a parity proof. The ray-basis symbol of this proof can be worked out as follows:
@@
In the first step we replaced each letter (or generator) by its profile as given in Table 6 and collected all pentadecagons occurring the same number of times over the profiles within brackets, using subscripts to indicate the number of times they occurred over the profiles. In the second step we replaced the string of letters within each pair of brackets by fifteen times the number of letters and added the number of bases (which is fifteen times the number of letters in the word) after the dash.
Table 8 shows examples of parity proofs of all lengths from one to fifteen obtained by combining the words in Table 7. The question arises as to which parity proofs of the 120-cell are minimal (i.e., cannot be reduced to shorter proofs by leaving out some of their letters). The answer can be given by using a result from coding theory, since the words describing parity proofs can be interpreted as the codewords of a linear binary code.
Recall that the matrix $M$ has dimensions $m\times n$, where $m$ is the number of pentadecagons and $n$ is the number of generators. The nullspace of $M$ has $k(<n)$ linearly independent vectors from which all the others can be built up. These $k$ vectors span the rowspace of an $(n,k)$ linear binary code whose odd length [^9] By an odd-length codeword we mean one having an odd number of 1’s in it. The length of a codeword is often termed its weight in coding theory. codewords correspond to fifteen-fold symmetric parity proofs. Thus the question of determining which parity proofs are minimal is the same as that of determining which odd length codewords of an $(n,k)$ linear binary code are minimal (i.e., cannot be converted into smaller odd length codewords by replacing some of their 1’s by 0’s). The answer to this question is partially given by the following result of Ashikhmin and BargBarg :
@@
Proposition 1. The weight $l$ of a minimal codeword in an $(n,k)$ linear binary code must satisfy $l\leq n-k+1$.
For the 120-cell, $n=45$ and $k=30$, so a minimal codeword must have length $l\leq 16$, implying that no parity proofs of length 17 or higher can be minimal. One can use the MacWilliams identities MacW1 ; MacW2 to calculate the number of parity proofs of any length and finds that the number of proofs of length 1,3,5,7 or 9 is 4,48,564,5116 or 42576, respectively. The number grows rapidly after that and peaks at well over a hundred million for proofs of length 23 before falling off at the high end (see Table 1 of Supplementary Materials for a complete list). The longest proofs have a length of 39 and there are 1212 of them. As mentioned above, no parity proof of length 17 or larger can be minimal. As for proofs of length 15 or less, a detailed examination is needed in each case to determine if it is minimal or not.
Fifteen-fold symmetric parity proofs, whether minimal or not, sometimes break up into a number of smaller parity proofs. We will term them reducible if they do, and irreducible if they do not. The smaller proofs within a reducible proof do not involve entire pentadecagons of rays and therefore cannot be described as words made up of the lower case letters we have introduced. We will describe them simply by their ray-basis symbols. Table 8 shows how all the reducible proofs break up into smaller proofs, with the direct sum symbol, $\oplus$, being used in those cases where the smaller proofs have no bases in common and the ordinary sum, +, being used in the cases they do. Table 9 shows a blown up version of the reducible proof cdycdy, which makes clear how its bases split up into those of three smaller proofs, each of which has a five-fold symmetry. The other reducible proofs in Table 8 also split up into a number of smaller proofs of five-fold symmetry, making them easy to pick out.
The smaller proofs in a reducible proof could have still smaller quantum contextual sets within them. These sets are simply sets of bases whose rays cannot be assigned the value 0 or 1 in a noncontextual fashion in such a way that every basis has exactly one ray assigned the value 1 in it. Only a detailed examination in each case can tell if this is so or not.

### v3

@@ A A D D B B C C A B C D A B C D A B C D
Table 4: Basis Table of the 600-cell. The generators $a$ through $e$ are shown in bold font at the tops of the columns, just below their basis profiles (also in bold font), and the orbit of each generator is shown below it. Any basis in an orbit is obtained from the one above it by adding one to each of its rays, but with “wraparound”. To understand what this means, consider the basis 25,27,45,37 in the second column. Adding one to each of its rays gives 26,28,46,38. However the third ray, 46, has overshot the boundary of the pentadecagon C to which its predecessor belonged, and so should be replaced by the ray 31 at the beginning of that pentadecagon. Several other instances of wraparound occur in the table. The 75 bases shown here make up the entire basis table of the 600-cell. The orbits of the generators $a$ and $b$ each give rise to a $30_{2}-15_{4}$ parity proof, but the orbits of the other generators do not.
Table 4 shows the orbits of fifteen bases arising from the generators $a$ through $e$ of Table 3. The 75 bases shown there make up the entire basis table of the 600-cell, which can be described by the word abcdeabcde. The bases associated with the generators $a$ and $b$ each yield a $30_{2}-15_{4}$ parity proof involving thirty rays that each occur twice over an odd number (fifteen) of bases.
One can ask if the 600-cell has any fifteen-fold proofs beyond the two found above. Any such proof must be expressible as a “word” made up of an odd number of distinct letters from $a,b,c,d$ and $e$ (to ensure that the total number of bases in it is odd). However, it must also have each of the pentadecagons A,B,C and D occur an even number of times over its letters to ensure that each ray occurs an even number of times in it. It is easy to see that these two conditions allow only the eight proofs described by the words $a,b,acd,ace,ade,bcd,bce$ and bdebde, of which the six three-letter words are the new ones. The ray-basis symbol associated with any three-letter proof can be obtained by replacing each of its letters/ generators by its basis profile, counting the number of times each of the pentadecagons occurs over the word and finally converting the pentadecagon counts into ray counts. We illustrate how this works for the word acdacd below:
acdacd $\rightarrow$ (AADD)$(ABCD)$ $\rightarrow$ (CD)2(AB)4 $\rightarrow$$30_{2}30_{4}-45_{4}$   (1)
All the fifteen-fold proofs of the 600-cell are shown in Table 6, along with their ray-basis symbols.
Parity proof Pentadecagon counts Ray-Basis symbol
$a$ (AD)2 302–154
$b$ (BC)2 302–154
acdacd, aceace, adeade (BC)2(AD)4 $30_{2}30_{4}-45_{4}$
bcdbcd, bcebce, bdebde (BC)2(AD)4 $30_{2}30_{4}-45_{4}$
Table 5: Fifteen-fold parity proofs of the 600-cell. The first column shows the word(s) representing the proof(s), the second its/their pentadecagon counts and the third the ray-basis symbol. All proofs with the same entries in the second and third columns have been lumped together in the first column.
We will term a parity proof irreducible if no subset of its bases constitutes a parity proof of any kind (even one of less than fifteen-fold symmetry) and reducible otherwise. Of the proofs in Table 5, only the single letter proofs are irreducible while the six three-letter proofs are reducible because they can all be reduced to one of the single letter proofs $a$ or $b$ by dropping two of their letters.
@@ 5 The 120-cell
The 120-cell is a four-dimensional regular polytope with 600 vertices lying on the surface of a 3-sphere. Its vertices come in antipodal pairs, so it has 300 rays. The rays form 675 bases, with each ray occurring in nine bases, so the ray-basis symbol of this polytope is $300_{9}-675_{4}$.
The Kochen-Specker diagram obtained from the triacontagonal projection arranges the rays at the vertices of twenty concentric regular pentadecagons, sixteen of which lie in pairs on eight different circles and four of which lie on circles of their own. Following ChiltonChilton , we will label the pentadecagons by the letters A through L in order of decreasing circumradius, but add subscripts to eight of the letters to distinguish between pentadecagons lying on circles of the same radius. The numbering scheme we use for the rays is shown in Table 6. The main difference from the 600-cell is that there are now pairs of pentadecagons on a common circumcircle, with the vertices of one interleaving those of the other. In such cases, we let the numbers increase all the way around one pentadecagon before we let them increase around the other (see Table 6 for details).
The 675 bases [^6] If one calculates the bases using triacontagonal coordinates, then the generators in Table 7 can be picked out very easily. of the 120-cell fall into forty five different orbits of fifteen bases each. Table 7 lists one generator for each orbit, along with its basis profile. Most generators have a unique profile, but there are two sets of three generators that each have a common profile. All the generators give rise to distinct orbits, irrespective of whether they share a profile with others or not.
Four of the generators in Table 7, namely, the ones labeled $j,q,r^{\prime}$ and $s^{\prime}$ give rise to $30_{2}-15_{4}$ parity proofs. However it is possible to obtain a large number of other proofs by combining the generators in Table 7 suitably, and we proceed to show how this can be done.
Pentadecagon Rays
A 1-15
B1 16-30
B2 31-45
C 46-60
D1 61-75
D2 76-90
E1 91-105
E2 106-120
F1 121-135
F2 136-150
G1 151-165
G2 166-180
H1 181-195
H2 196-210
I1 211-225
I2 226-240
J 241-255
K1 256-270
K2 271-285
L 286-300
Table 6: Pentadecagons of the 120-cell, labeled A through L in order of decreasing circumradius, with pairs having the same circumradius being distinguished from each other by a subscript. The rays are numbered in counter-clockwise order around the pentadecagons, beginning from the outermost one and proceeding inwards and with the lowest number on any pentadecagon being attached to the ray lying on or just above the positive x-axis. For pentadecagons lying on a common circumcircle, the numbers first increase around the one with the smaller subscript before they increase around the other.
@@
Table 7: Generators of the 120-cell. The generators, which are labeled by lower case letters, are shown in the second and fourth columns, with the basis of each shown in curly brackets. The basis profile of each generator is shown to its left, with each capital letter indicating the pentadecagon to which the corresponding ray of the generator belongs. The generators $n$,$o$ and $p$ all have a common profile, as do $u$,$v$ and $w$. Thus there are just 41 different profiles for the 45 generators. Each generator gives rise to an orbit of fifteen bases and the union of all the orbits gives the 675 bases of the 120-cell.
We use a technique due to Lisonĕk et allisonek1 , suitably adapted to our needs, to solve the problem. Consider the matrix $M$ whose $i,j$-th element, $M_{i,j}$, is the number of times pentadecagon $i$ of Table 6 occurs in the basis profile of generator $j$ of Table 7 (we number the pentadecagons and generators in the order they occur in these tables) [^7] It is here that our approach deviates from lisonek1 , which uses rays and bases in place of our pentadecagons and generators in constructing the matrix $M$. Thus our approach might be considered a coarse-grained version of their fine-grained one. While our approach is unable to capture all the parity proofs in the 120-cell, it compensates by offering a far more detailed view of the ones it does.. The matrix $M$ has dimensions 20 x 45. A parity proof can be extracted from $M$ by picking an odd number of its columns in such a way that if they are stacked side by side, the sum of the numbers in each of the horizontal rows is even. The word describing the resulting proof has for its letters the generators corresponding to the columns that are picked. It remains only to explain how the right columns of $M$ are to be picked.
The problem can be solved by finding all 45-dimensional column vectors $X$, with elements of 0 or 1, for which the equation $MX=0$ (mod 2) is satisfied, for then the non-vanishing elements of $X$ will indicate the columns of $M$ to be picked. The solutions to the preceding equation are all vectors in the nullspace of the matrix $M$ mod 2. We used Maple to find that the nullspace has dimension 30 and found a set of linearly independent vectors in it. Taking all possible linear combinations of these vectors, with coefficients of 0 or 1, gives $2^{30}$ solutions to the equation $MX=0$ (mod 2). However only half these solutions (namely, the ones with an odd number of 1’s as elements) give parity proofs. Thus the number of parity proofs is $2^{29}$, and each of them can be constructed as the union of the bases associated with the generators picked out by the non-vanishing elements of $X$.
@@
Table 8: The 30 words representing a set of linearly independent vectors in the nullspace of the matrix $M$. The first word is of odd length, while all the others are of even length. The letters of any word are always written in alphabetical order, with the primed letters written after the unprimed ones.
@@ 13 $bcdefghiknszq^{\prime}$ $225_{2}30_{4}15_{6}15_{8}$-$195_{4}$ (irreducible)
Table 9: Fifteen-fold proofs of the 120-cell of all word lengths from 1 to 15 obtained by combining words in Table 8 according to the rule in Eq.(2). Words/proofs with the same ray-basis symbol have been lumped together in the first column. If the proof is reducible, the way its ray-basis symbol splits up into those of the smaller proofs in it is indicated in the last column (see the text for an explanation). The symbol $\oplus$ is used if the smaller proofs have no bases in common, whereas + is used if they do (this happens only for the last proof). See Table 10 for a blown-up version of the proof cdycdy, which shows how its bases split up into those of the three smaller proofs in it. Note: the dots in the last column indicate that the ray-basis symbol after the equals sign is to be added to itself the number of times indicated in brackets at the end, with the type of addition, $\oplus$ or +, being the same as before.
@@ A C J L A D1 I1 L C D1 I1 J
Table 10: The three-letter proof cdycdy. The generators $c$,$d$ and $y$ are shown in bold font at the tops of the three columns, just below their basis profiles, and their orbits of fifteen bases are shown below them. The 45 bases in this table give a $90_{2}-45_{4}$ parity proof, but this proof is reducible and breaks up into the three smaller $30_{2}-15_{4}$ proofs shown in bold, italic and regular font. Each of the smaller proofs has a five-fold symmetry (in that adding three to the rays of any basis yields the next basis in the chain). The smaller proofs have no bases in common, so cdycdy is written as a direct sum ($\oplus)$ of them in Table 9.
The 30 linearly independent vectors in the nullspace of $M$ can each be written as a “word” made up of distinct letters representing the generators in Table 7. One choice for these words is shown in Table 8, with the words listed in order of increasing length and the letters of any word always written in alphabetical order, with the primed letters coming after the unprimed ones (this convention eliminates the confusion caused by encountering the same word written in different forms). Henceforth, by the “length” of a fifteen-fold proof, we will mean the number of letters in the word representing it.
We now give a simple algorithm for obtaining all the fifteen-fold parity proofs. It consists of the following two steps: (1) pick an arbitrary subset of the 29 even length words in Table 8 and combine them together in the manner explained in the next paragraph to obtain an even length word, and (2) combine this word with the letter $j$ to obtain an odd length word representing a parity proof. The bases of the proof can be obtained as the union of the orbits of the letters/generators in its word.
The rule for combining two words alluded to above is the following. Let $U$ and $V$ be the sets of letters in the two words. Then the word obtained by combining them is represented by the set $W$ that is the symmetric difference of the sets $U$ and $V$:
@@
Stated in words (pun unintended!), Eq.(2) says that the word obtained by combining two words is the union of their sets of letters, but with all letters common to the two words dropped. We will actually use this rule to combine any two words, whether they belong to Table 8 or not.
The first step of our algorithm can be executed with the aid of Eq.(2) as follows: one combines any two words from the chosen subset of Table 8 to obtain a new word and then combines it with one of the remaining words from the subset, and keeps proceeding in this way until all the words in the subset get used up. The new words that result at each step of this process all lie outside the list in Table 8, with the final (even length) word being the input for the second step of the algorithm.
An example should help make the above algorithm clear. Suppose one picks just the two even length words abdekrabdekr and $dgji^{\prime}$ from Table 8. Then, in the first step of the algorithm, one combines these words to get the 8-letter word $abegjkri^{\prime}$, and in the second step one combines it with $j$ to get the 7-letter word $abegkri^{\prime}$ that represents a parity proof. The ray-basis symbol of this proof can be worked out as follows:
@@
The transition from the word to its ray-basis symbol can be made in the same way as for the 600-cell in Eq.(1) by using the basis profiles in Table 7.
The algorithm we have just given for obtaining fifteen-fold proofs can be rephrased in terms of adding vectors in GF($2^{45}$). To do this, we first write each of the vectors in the nullspace of $M$ as a string of 45 bits whose 1’s represent the letters/generators present in the corresponding word. The operation of combining a subset of the words in Table 8 can then be carried out by adding the corresponding 45-bit strings bitwise modulo 2 and converting the resulting string back into the word it represents. If the second step of the algorithm is omitted, only even length words are obtained, but adding the second step converts the even words into an equal number of odd length words that span an affine subspace of the same dimension as the even words.
Table 9 shows examples of proofs of word lengths one to fifteen obtained by combining the words in Table 8, with the last column indicating if the proof is irreducible or not. For each reducible proof, the smaller proofs in it are indicated by their ray-basis symbols, with the $\oplus$ sign being used if the proofs have no bases in common and the + sign if they do. One can determine if a proof is irreducible or not by setting up a matrix $M$ like the one we used earlier, but between the individual rays and bases of the proof, and determining the dimension of its nullspace; if the dimension is 1 the proof is irreducible, whereas if it larger than 1 it is reducible and one can determine the bases of all the smaller proofs in it. Table 10 shows how the reducible proof cdycdy of Table 9 splits up into three smaller proofs that have no bases in common (hence the $\oplus$ sign). Note also that each of the smaller proofs has a five-fold symmetry, as explained in more detail in the caption to the table.
The proofs in Table 9 are just a vanishingly small fraction of all the fifteen-fold proofs that exist. To get a better idea of the proofs as a whole, we note that because each proof can be written as a 45-bit string of 1’s and 0’s, the proofs are like the codewords of odd weight [^8] The weight of a binary codeword is the number of 1’s in it. of a linear binary code [45,30,1][45,30,1] [^9] An $[n,k,d]$ linear binary code consists of binary codewords of $n$ bits of which $k$ are used to transmit a message and the other nkn-k are used to correct errors and $d$ is the minimum distance between codewords (the distance between two codewords is the number of places in which their bits differ). spanned by the 30 codewords in Table 8. One can use the MacWilliams identities MacW1 ; MacW2 to calculate the number of codewords of each odd weight (and hence the number of parity proofs of each odd word length) and finds that the number of proofs of word length 1,3,5,7 or 9 is 4,48,564,5116 or 42576, respectively. The number grows rapidly after that and peaks at well over a hundred million for proofs of word length 23 before falling off to zero after a length of 39 (see Table 1 of Supplementary Materials for a complete histogram of the codewords).
How many of the above proofs are reducible? While it is possible to tell by a direct examination if any given proof is reducible or not, is there a general statement about reducibility that one can make about a large class of proofs without having to examine every one? It turns out there is. Recall that a reducible proof is one whose word can be made to pass into the word of a smaller proof by leaving out an even number of the letters in it (such proofs of the 600-cell were exhibited in Table 5). In coding theory, a binary codeword is said to be minimal if it cannot be made to pass into another codeword by replacing some of its 1’s by 0’s. The following result of Ashikhmin and BargBarg provides some information about which codewords might be minimal:
@@
Proposition 1. The weight $l$ of a minimal codeword in an $[n,k,d]$ linear binary code must satisfy the inequality $l\leq n-k+1$.
For the code associated with the 120-cell, $n=45$ and $k=30$ and Proposition 1 implies that a minimal codeword must have weight $l\leq 16$. Translating this into the language of our proofs, we can conclude, at one stroke, that no fifteen-fold proof of word length 17 or larger can be irreducible. However the proposition is ambivalent about proofs of length 15 or less and only a detailed examination will tell if any particular proof is irreducible or not. We have actually carried out such an examination of all the proofs in Table 9 and listed the results in the last column of that table.
To sum up, while Proposition 1 implies that fifteen-fold proofs of greater than a certain length are necessarily reducible, it says nothing about proofs of shorter length.